OpenStudy (anonymous):

consider the following quadratic equation: 49x^2=25 using the standard form ax^2+bx+c=0 of the quadratic equation above, factor the left hand side equation into two linear factors

OpenStudy (anonymous):

49x^2=25 49x^2-25=0 (7 x-5) (7 x+5)=0